OpenStudy (anonymous):

Find a cubic function with the given zeros. You must show your work for credit. -6, 5, -2 for a medal please

OpenStudy (australopithecus):

(x + 6)(x - 5)(x + 2)

OpenStudy (australopithecus):

easy

OpenStudy (anonymous):

Thats it ?

OpenStudy (australopithecus):

yup

OpenStudy (anonymous):

(x+6)(x-5)(x+2)

OpenStudy (australopithecus):

it is a cubic function that is zero at all the points listed

OpenStudy (anonymous):

Yess ,

OpenStudy (anonymous):

Is there any work to that i need it ?

OpenStudy (anonymous):

No

OpenStudy (australopithecus):

If you want you can multiply it out so it isn't in factored form

OpenStudy (anonymous):

Yeaa ,

OpenStudy (anonymous):

sure can you put please

OpenStudy (australopithecus):

You do not have to do that though

OpenStudy (australopithecus):

Although if you do not know how to do that then you should probably learn

OpenStudy (australopithecus):

especially if you are in functions

OpenStudy (anonymous):

Right You Can Just Do The First Part ,

OpenStudy (anonymous):

so I would say for the cubic function with the given zeros you would have the given points (x+6)(x-5)(x+2) That would be my answer?

OpenStudy (anonymous):

Right

OpenStudy (australopithecus):

Do you not understand how we got that answeR?

OpenStudy (anonymous):

not really explain

OpenStudy (anonymous):

(a+b)(c+d)(e+f) = ( a(c) + a(d) + b(c) + b(d) )(e+f) )

OpenStudy (australopithecus):

(a+b)(c+d)(e+f) = ( a(c) + a(d) + b(c) + b(d) )(e+f) )

OpenStudy (anonymous):

Do Youu Get It

OpenStudy (australopithecus):

We know (x + 6) when x = -6 (-6 + 6) = 0 (x -5) when x = 5 (5 - 5) = 0 etc we know that a cubic function is a function that has x^3 thus if we multiply all three together we will get an x^3

OpenStudy (australopithecus):

This will give x^2 (x+3)(x+2) This will give x^3 (x + 6)(x - 5)(x + 2) this's will give x^4 (x+3)(x+4)(x+4)(x+4)

OpenStudy (australopithecus):

etc

OpenStudy (australopithecus):

When the only stipulation is that a function has an x to a specific degree and particular zeros you can just make up a factored form that has these traits vary easily

OpenStudy (anonymous):

so hows this : for the cubic function with the given zeros you would have the given points (x+6)(x-5)(x+2)(x + 6) for a further example I know that x = -6 because(-6+6)=0 and also(x -5) x = 5 (5-5)=0 and furthermore you know that a this is a cubic function if it has x^3 and if I were to multiply them all I would get x^3 (x + 6)(x - 5)(x + 2)= x^3

OpenStudy (anonymous):

Thats Correct

OpenStudy (australopithecus):

(x + 6)(x - 5)(x + 2) = (x^2 - 5x + 6x + 30)(x+2) = (x^2 + x + 30)(x+2) = x^3 + 2x^2 + x^2 + 2x + 30x + 60 = x^3 + 3x^2 + 32x + 60

OpenStudy (anonymous):

Right ^